# VGGT-SLAM 2.0: Real time Dense Feed-forward Scene Reconstruction

### v1

@@ Abstract
We present VGGT-SLAM 2.0, a real time RGB feed-forward SLAM system which substantially improves upon VGGT-SLAM for incrementally aligning submaps created from VGGT. Firstly, we remove high-dimensional 15-degree-of-freedom drift and planar degeneracy from VGGT-SLAM by creating a new factor graph design while still addressing the reconstruction ambiguity of VGGT given unknown camera intrinsics. Secondly, by studying the attention layers of VGGT, we show that one of the layers is well suited to assist in image retrieval verification for free without additional training, which enables both rejecting false positive matches and allows for completing more loop closures. Finally, we conduct a suite of experiments which includes showing VGGT-SLAM 2.0 can easily be adapted for open-set object detection and demonstrating real time performance while running online onboard a ground robot using a Jetson Thor. We also test in environments ranging from cluttered indoor apartments and office scenes to a 4,200 square foot barn, and we also demonstrate VGGT-SLAM 2.0 achieves the highest accuracy on the TUM dataset with about 23 percent less pose error than VGGT-SLAM. Code will be released upon publication.
Figure 1: A VGGT-SLAM 2.0 map of an apartment scene showing a living room, kitchen, bedroom, 2 bathrooms, and closet using uncalibrated RGB images collected with an iPhone. After the map is constructed, it is easy to use it for retrieval of open-set objects. Here, eight example open-set object queries are included showing the resulting 3D bounding box produced by VGGT-SLAM 2.0 given the provided text query.
@@
This work was supported in part by the NSF Graduate Research Fellowship Program under Grant 2141064, and the ONR RAPID program.
@@ I Introduction
Recently, the foundational task in robotics and computer vision of using images from a camera to simultaneously create a 3D reconstruction of a scene and localize the camera has seen a paradigm shift from using fully classical multi-view geometry and optimization techniques [1] towards building on top of feed-forward geometric foundation models [2, 3, 4] to develop SLAM systems [5, 6, 7, 8, 9]. These new hybrid SLAM systems, which combine both geometric foundation models with tools from classical SLAM, produce a much simpler SLAM system that is both easier to use and maintain. They also produce dense RGB point clouds maps while not requiring camera calibration be known beforehand.
One recent work, VGGT-SLAM [5], shows that even though VGGT [4] is unable to directly process potentially thousands of frames as typically done in SLAM [^1] VGGT is limited to about 60 frames on an RTX 4090 with 24 GB due to memory constraints, it can be extended to become a SLAM system by creating and aligning submaps from VGGT where neighboring submaps share an overlapping frame. The overlapping frame is a common image between submaps which is used for alignment. Furthermore, VGGT-SLAM shows that simply aligning the submaps with a similarity transformation (rotation, translation, and scale) is not always sufficient since the VGGT submaps — which rely on uncalibrated camera images — sometimes have a higher dimensional projective ambiguity. VGGT-SLAM thus optimizes a 15 degree-of-freedom (DoF) alignment on the $\mathrm{SL}(4)$ manifold which compensates for submap distortions.
@@
However, there are several key limitations which hinder the deployment of VGGT-SLAM which are addressed in this paper. The first is the higher dimensional 15 DoF alignment introduces rapid drift between loop closures which can severely warp a scene before loop closures (Fig. 2) and is not always recoverable even with loop closures. Solving for the 15 DoF homography is also degenerate in planar scenes which can cause divergence in common scenarios such as a camera looking only at a wall or at a flat floor.
Secondly, the factor graph in VGGT-SLAM only attempts to estimate the homography between submaps instead of correcting keyframe level errors. This means that any potential translation and rotation error from VGGT is handled sub-optimally during global optimization from loop closures.
@@
  • Our second contribution (Sections IV-A and IV-C) is a new factor graph structure where all keyframes are nodes and the graph contains intra edges connecting keyframes inside a submap and inner edges connecting overlapping frames between submaps.
@@
  • Finally, to show the applicability of VGGT-SLAM we conduct a suite of experiments showing: (1) simple yet effective integration of open-set object detection into the geometric VGGT-SLAM map, (2) real time mapping on a Jetson Thor running online on board a ground robot, (3) environments ranging from cluttered indoor scenes to a 4,200 square foot barn to outdoor driving sequences from the Kitti dataset, and (4) the best accuracy of recent learning-based methods on the TUM RGB-D dataset with a 23 percent error decrease from VGGT-SLAM.
@@ II Related Works
Feed-forward scene reconstruction. Recently, following the seminal works of DUSt3R [2] and MASt3R [3] which are trained to take in a pair of uncalibrated monocular images and produce a 3D scene reconstruction, there has been a paradigm shift towards using Geometric foundation models (GFM) for SLAM. Since the two aforementioned works can only process a pair of images, multiple works have begun developing approaches to process larger numbers of frames. MASt3R-SFM [24] builds on MASt3R to perform global optimization over multiple images but quickly grows computationally expensive as the number of frames increases. MASt3R-SLAM [6] is the first work to demonstrate real-time SLAM with a GFM. MASt3R-Fusion [9] incorporates IMU and GNSS as additional sensor measurements through classical optimization.
Several works have developed GFMs to directly process multiple images [25, 26, 27, 4] with some also modifying the networks to also enable Gaussian Splatting [28] reconstructions [29]. However, all of these works are restricted by GPU memory usage to a relatively small number of frames. To extend VGGT [4] to a large number of frames, VGGT-SLAM [5] is the first work to incrementally create and align submaps with VGGT. VGGT-Long [8] aligns VGGT submaps for large-scale outdoor self-driving scenes. Recently, SING3R-SLAM [30] creates submaps and fuses them into a global Gaussian Splatting map and ViSTA-SLAM [7] uses a lightweight model for two-view association combined with a $\mathrm{Sim}(3)$ pose graph. Additionally, MegaSAM [31] builds upon GFMs to handle reconstruction for dynamic scenes and TTT3R [32] uses Test-Time Training to extend CUT3R [25] for longer sequences.
@@
III Notation and Prelimaries
Notation. We represent matricies with bold uppercase letters (e.g., C), sets using capital calligraphic fonts (e.g., $\mathcal{C}$), and vectors with bold lowercase letters (e.g., c). To create a submap, $\mathcal{S}_{i}$, we pass a set of frames $\mathcal{I}$ to VGGT, and use the output camera calibrations, $\mathcal{K}$; poses, $\mathcal{T}$; depth maps, $\mathcal{D}$; and corresponding depth confidence maps, $\mathcal{C}$. All submaps have $n$ keyframes, except for the loop closure submaps (Section IV-C) which have 2 frames.
Prelimaries. VGGT estimates 3D points which are defined w.r.t. the first camera of a submap. 3D points can be computed from either the point maps or by transforming the estimated depth by estimates of the calibrations and poses. Here, we will instead define 3D points w.r.t. their respective camera frame. Thus, the points X for frame $i$ (i.e., $\textbf{X}_{i}$) are computed through back projection using only $\textbf{K}^{-1}_{i}$ and $\textbf{D}_{i}$, where $\textbf{X}_{i}\in~{{\mathbb{R}}}^{3\times h\times w}$ for image of size $h\times w$.
@@ III Notation and Prelimaries
where we use overloaded notation such that X is in homogenous coordinates when multiplied by a homography. The $4\times 4$ homography matrix has 15 degrees of freedom: 3 for translation, 3 for rotation, 1 for scale, 5 for affine (i.e., camera calibration), and 3 for projective components [38]. Since the homography matrix is only up to scale, optimizing the residuals between homography matrices can be done by mapping each homography to the $\mathrm{SL}(4)$ manifold which consist of all $4\times 4$ matrices with unit determinant. This is a unique mapping and allows for factor graph optimization on the $\mathrm{SL}(4)$ manifold [5]. More common manifolds used in robotics such as $\mathrm{SE}(3)$ and $\mathrm{Sim}(3)$ are subsets of $\mathrm{SL}(4)$.
@@
Our objective is to incrementally align submaps produced by VGGT into a globally consistent map that recovers the true geometry of the scene (up to a similarity transform ambiguity as we do not estimate the true scene scale). To do this, in Section IV-A we create a factor graph where every node corresponds to a keyframe in a submap and edges describe the relationship between keyframes. Consecutive submaps share an image such that the first image of a new submap is the same as the last image of the prior submap. We refer to the two submaps’ respective estimate of the frame of this image as the overlapping frames. In Section IV-B we show how the attention layers of VGGT can be used for image retrieval verification, that is, given a queried frame and a retrieved frame which are predicted to have overlap by a third party image retrieval method (in our case SALAD [10]), we can use the attention layers to provide assurance that VGGT identifies overlap in the images. Finally, in Section IV-C we show how our new factor graph structure and image retrieval verification are combined to enable global optimization with loop closures.
Figure 3: Factor graph structure of VGGT-SLAM 2.0 showing edges inside and between submaps where all keyframes are nodes and loop closure submaps consist of two frames. Nodes connected with an inner submap edge such as $H_{3}$ and $H_{4}$ are an example of overlapping nodes (meaning the camera image associated with these nodes is identical). Here, for the loop closure submap which is made up of nodes $H_{14}$ and $H_{15}$, $H_{12}$ represents the node of a queried frame which was matched with the node $H_{3}$ of the retrieved frame.
@@ IV-A Relative frame alignment
For edges inside a submap (intra edges), the edges will initially only have non-identity rotation and translation (i.e., $\mathrm{SE}(3)$) variables since by construction, the relationship of 3D points, $X_{i}$ for each frame inside a single run of VGGT can be expressed using Eq. 1 where the homography only has $\mathrm{SE}(3)$ components. Optimizing these edges will help correct potential drift in VGGT’s translation and rotation estimates. For edges between overlapping frames (inner edges) there will only be non-identity calibration (i.e., affine) and scale variables which will be used to make each submap’s estimate of the scale and calibration of the overlapping frame be identical. Since both of these transformations are subsets of $\mathrm{SL}(4)$, we use $\mathrm{SL}(4)$ factors for factor graph optimization. A visualization of the factor graph structure showing different edge types can be seen in Fig. 3. Importantly, given that there is error in VGGT’s estimate of calibration, we do not know what the true camera calibration is for the overlapping frames — only that they must have the same calibration.
Intra submap frame alignment. For two keyframes $i$ and $j$ inside a submap, the homography $\mathbf{H}^{i}_{j}$ from Eq. 1 that transfers correspondences between point clouds $\mathbf{X}^{i}$ and $\mathbf{X}^{j}$ is simply a rotation and translation which we can get directly from the VGGT outputted poses as:
@@
Inner submap frame alignment. Now, we have to find the alignment between the overlapping frames $i$ and $j$ in overlapping submaps. For example, nodes $H_{3}$ and $H_{4}$ in Fig. 3 correspond to overlapping frames. Importantly, instead of solving for a 15 DoF transformation by optimizing an alignment over the 3D points corresponding to the two overlapping frames as was done in VGGT-SLAM, here we recognize that the overlapping frames by construction must have the same translation, rotation, and camera calibration. Thus, we enforce that the relative translation and rotation between the overlapping frames is zero and determine the transformation needed to align their calibration. We then solve for the scale factor, $s$.
The homography is then:
@@
where the calibration component of the homography comes from the VGGT estimated calibrations. Importantly, even if the calibrations $K_{i}$ and $K_{j}$ do not represent the true camera calibration (which can occur if VGGT incorrectly estimates camera calibration which results in an affine warping of a submap), this enforces their calibrations to be identical. We know they must be identical because the overlapping frames represent the same camera. Solving for the scale factor is the only time we directly use the VGGT 3D points. To find $s$, we leverage two ideas. The first is that since the overlapping frames are identical, corresponding 3D points between them are trivially known. The second is that unlike in [5], with our new problem formulation, the scale factor estimate is invariant to all other parameters of the homography and to the VGGT estimated poses. We are invariant to $\mathrm{SE}(3)$ components since the 3D points of each image are defined not w.r.t. the first frame of a submap but rather to their own camera frame. We make the scale estimate invariant to the affine components by first warping the point clouds $\mathbf{X}^{i}$ and $\mathbf{X}^{j}$ to have the same calibration using $\textbf{K}_{i}^{-1}$ and $\textbf{K}_{j}$ and then estimate the scale factor as the median ratio between corresponding 3D point distances for all points with sufficient depth confidence, $\mathcal{C}$.
@@ IV-B Image retrieval verification with VGGT
In this section, we demonstrate that by studying the attention layers of VGGT, we can easily extract information to determine if VGGT is able to find overlap between input images. While we find our analysis holds for an arbitrary number of frames passed to VGGT, since our objective in this section is image retrieval verification, we focus on the scenario where VGGT is given two images (a query and a retrieved frame). In particular, we reveal that layer 22 in VGGT consistently pinpoints whether VGGT predicts correspondences between the two images. In Fig. 4 we visualize the attention matrix of a selected key token (shown with a black star), with respect to all other query tokens. Layer 22 consistently shows a spotlight like attention map which shows larger attention values at the correspond location of the image. This is not clearly present in the other layers, including the nearby layers (21 and 23) which are shown in Fig. 4 for comparison. This spotlight affect also works on low texture regions of a scene as demonstrated by the first example of Fig. 4 which places the query token on a plain white wall.
@@
(a) Queried and retrieved frames that have overlap. Our estimated match score: 1.026
(b) Queried and retrieved frames that do not have overlap. Our estimated match score: 0.491
(c) Queried and retrieved frames that do not have overlap, but SALAD incorrectly estimated a match between them. Our estimated match score: 0.550

### v2

# VGGT-SLAM 2.0: Real-time Dense Feed-forward Scene Reconstruction

@@ Abstract
We present VGGT-SLAM 2.0, a real-time RGB feed-forward SLAM system which substantially improves upon VGGT-SLAM for incrementally aligning submaps created from VGGT. Firstly, we remove high-dimensional 15-degree-of-freedom drift and planar degeneracy from VGGT-SLAM by creating a new factor graph design while still addressing the reconstruction ambiguity of VGGT given unknown camera intrinsics. Secondly, by studying the attention layers of VGGT, we show that one of the layers is well suited to assist in image retrieval verification for free without additional training, which enables both rejecting false positive matches and allows for completing more loop closures. Finally, we conduct a suite of experiments which includes showing VGGT-SLAM 2.0 can easily be adapted for open-set object detection and demonstrating real-time performance while running online onboard a ground robot using a Jetson Thor. We test in environments ranging from cluttered indoor apartments and office scenes to a 4,200 square foot barn, and we also demonstrate VGGT-SLAM 2.0 achieves the highest accuracy on the TUM dataset with about 23 percent less pose error than VGGT-SLAM. Code will be released upon publication.
Figure 1: A VGGT-SLAM 2.0 map of an apartment scene including a living room, kitchen, bedroom, 2 bathrooms, and closet built using uncalibrated RGB images collected with an iPhone. After the map is constructed, it is easy to use it for retrieval of open-set objects. Here, eight example open-set object queries are included, showing the resulting 3D bounding box produced by VGGT-SLAM 2.0 given the provided text query.
@@
This work was supported in part by the NSF Graduate Research Fellowship Program under Grant 2141064, the ARL DCIST program, and the ONR RAPID program.
@@ I Introduction
Recently, the foundational task in robotics and computer vision of using images from a camera to simultaneously create a 3D reconstruction of a scene and localize the camera has seen a paradigm shift from using classical multi-view geometry and optimization techniques [1] towards building on top of feed-forward geometric foundation models [2, 3, 4] to develop SLAM systems [5, 6, 7, 8, 9]. These new hybrid SLAM systems, which combine both geometric foundation models with tools from classical SLAM, produce a much simpler SLAM system that is both easier to use and maintain. They also produce dense RGB point clouds maps while not requiring camera calibration be known beforehand.
One recent work, VGGT-SLAM [5], shows that even though VGGT [4] is unable to directly process potentially thousands of frames as typically done in SLAM [^1] VGGT is limited to about 60 frames on an RTX 4090 with 24 GB due to memory constraints., it can be extended to become a SLAM system by creating and aligning submaps from VGGT where neighboring submaps share an overlapping frame. The overlapping frame is a common image between submaps which is used for alignment. Furthermore, VGGT-SLAM shows that simply aligning the submaps with a similarity transformation (rotation, translation, and scale) is not always sufficient since the VGGT submaps — which rely on uncalibrated camera images — sometimes have a higher dimensional projective ambiguity. VGGT-SLAM thus optimizes a 15-degree-of-freedom (DoF) alignment on the $\mathrm{SL}(4)$ manifold which compensates for submap distortions.
@@
However, there are several key limitations which hinder the deployment of VGGT-SLAM which are addressed in this paper. The first is the higher dimensional 15-DoF alignment introduces rapid drift between loop closures which can severely warp a scene before loop closures (Fig. 2) and is not always recoverable even with loop closures. Solving for the 15-DoF homography is also degenerate in planar scenes which can cause divergence in common scenarios such as a camera looking only at a wall or at a flat floor.
Secondly, the factor graph in VGGT-SLAM only attempts to estimate the homography between submaps instead of correcting keyframe-level errors. This means that any potential translation and rotation error from VGGT is handled sub-optimally during global optimization from loop closures.
@@
  • Our second contribution (Sections IV-A and IV-C) is a new factor graph structure where all keyframes are nodes and the graph contains intra edges connecting keyframes inside a submap and inter edges connecting overlapping frames between submaps.
@@
  • Finally, to show the applicability of VGGT-SLAM we conduct a suite of experiments showing: (1) simple yet effective integration of open-set object detection into the geometric VGGT-SLAM map, (2) real-time mapping on a Jetson Thor running online on board a ground robot, (3) reconstruction of environments ranging from cluttered indoor scenes to a 4,200 square foot barn to outdoor driving sequences from the Kitti dataset, and (4) the best accuracy of recent learning-based methods on the TUM RGB-D dataset with a 23% error decrease from VGGT-SLAM.
@@ II Related Works
Feed-forward scene reconstruction. Recently, following the seminal works of DUSt3R [2] and MASt3R [3], which are trained to take in a pair of uncalibrated monocular images and produce a 3D scene reconstruction, there has been a paradigm shift towards using Geometric foundation models (GFM) for SLAM. Since the two aforementioned works can only process a pair of images, multiple works have begun developing approaches to process larger numbers of frames. MASt3R-SFM [24] builds on MASt3R to perform global optimization over multiple images but quickly grows computationally expensive as the number of frames increases. MASt3R-SLAM [6] is the first work to demonstrate real-time SLAM with a GFM. MASt3R-Fusion [9] incorporates IMU and GNSS as additional sensor measurements through classical optimization.
Several works have developed GFMs to directly process multiple images [25, 26, 27, 4] with some also modifying the networks to also enable Gaussian Splatting [28] reconstructions [29]. However, all of these works are restricted by GPU memory usage to a relatively small number of frames. To extend VGGT [4] to a large number of frames, VGGT-SLAM [5] is the first work to incrementally create and align submaps with VGGT. VGGT-Long [8] aligns VGGT submaps for large-scale outdoor self-driving scenes. Recently, SING3R-SLAM [30] creates submaps and fuses them into a global Gaussian Splatting map and ViSTA-SLAM [7] uses a lightweight model for two-view association combined with a $\mathrm{Sim}(3)$ pose graph. Additionally, MegaSaM [31] uses deep visual SLAM to handle reconstruction for dynamic scenes and TTT3R [32] uses Test-Time Training to extend CUT3R [25] to longer sequences. Concurrently, TALO [33] uses a sparse set of control points and Thin Plate Spline deformations fields to align submaps created by a GFM.
@@
III Notation and Preliminaries
Notation. We represent matrices and vectors with bold letters (e.g., C) and sets using capital calligraphic fonts (e.g., $\mathcal{C}$). To create a submap, $\mathcal{S}_{i}$, we pass a set of frames $\mathcal{I}$ to VGGT, and use the output camera calibrations, $\mathcal{K}$; poses, $\mathcal{T}$; depth maps, $\mathcal{D}$; and corresponding depth confidence maps, $\mathcal{C}$. All submaps have $n$ keyframes, except for the loop closure submaps (Section IV-C) which have 2 frames.
Preliminaries. VGGT estimates 3D points which are defined w.r.t. the first camera of a submap. 3D points can be computed from either the point maps or by transforming the estimated depth by estimates of the calibrations and poses. In this paper, we will instead define 3D points w.r.t. their respective camera frame. Thus, the points X for frame $i$ (i.e., $\textbf{X}_{i}$) are computed through back projection using only $\textbf{K}^{-1}_{i}$ and $\textbf{D}_{i}$, where $\textbf{X}_{i}\in~{{\mathbb{R}}}^{3\times h\times w}$ for image of size $h\times w$.
@@ III Notation and Preliminaries
where we use overloaded notation such that X is in homogenous coordinates when multiplied by a homography. The full $4\times 4$ homography matrix,
$\textbf{H}_{i}=\begin{bmatrix}{\textbf{KR}}&\textbf{t}\\ \mathbf{v}^{T}&s\end{bmatrix}$, (2)
has 15 degrees of freedom: 3 for translation (t), 3 for rotation (R), 1 for scale (s), 5 for affine (i.e., camera calibration, K), and 3 for projective components (v[39]. Since the homography matrix is only up to scale, optimizing the residuals between homography matrices can be done by mapping each homography to the $\mathrm{SL}(4)$ manifold which consist of all $4\times 4$ matrices with unit determinant, allowing for factor graph optimization on the $\mathrm{SL}(4)$ manifold [5]. More common manifolds used in robotics such as $\mathrm{SE}(3)$ and $\mathrm{Sim}(3)$ are subsets of $\mathrm{SL}(4)$.
@@
Our objective is to incrementally align submaps produced by VGGT into a globally consistent map that recovers the true geometry of the scene (up to a similarity transform ambiguity as we do not estimate the true scene scale). To do this, in Section IV-A we create a factor graph where every node corresponds to a keyframe in a submap and edges describe the relationship between keyframes. Consecutive submaps share an image such that the first image of a new submap is the same as the last image of the prior submap. We refer to the two submaps’ respective estimate of the frame of this image as the overlapping frames. In Section IV-B we show how the attention layers of VGGT can be used for image retrieval verification, that is, given a queried frame and a retrieved frame which are predicted to have overlap by a third-party image retrieval method (in our case SALAD [10]), we can use the attention layers to provide assurance that VGGT identifies overlap in the images. Finally, in Section IV-C we show how our new factor graph structure and image retrieval verification are combined to enable global optimization with loop closures.
Figure 3: Factor graph structure of VGGT-SLAM 2.0 showing edges inside and between submaps where all keyframes are nodes and loop closure submaps consist of two frames. Nodes connected with an inter submap edge such as $H_{3}$ and $H_{4}$ are an example of overlapping nodes (meaning the camera image associated with these nodes is identical). Here, for the loop closure submap which is made up of nodes $H_{14}$ and $H_{15}$, $H_{12}$ represents the node of a queried frame which was matched with the node $H_{3}$ of the retrieved frame.
@@ IV-A Relative frame alignment
For each keyframe, we want to estimate a homography $\textbf{H}_{i}$ which transforms the VGGT reconstruction (which may have submap-specific distortions) from the local camera frame to a globally consistent map. For edges inside a submap (intra edges), the edges will only have non-identity rotation and translation (i.e., $\mathrm{SE}(3)$) variables since projective distortion is consistent inside a submap, which means we do not have to apply affine correction to each keyframe separately. Optimizing these edges will help correct potential drift in VGGT’s translation and rotation estimates. For edges between overlapping frames (inter edges) there will only be non-identity calibration (i.e., affine) and scale variables which will be used to make each submap’s estimate of the scale and calibration of the overlapping frame be identical. Since both of these transformations are subsets of $\mathrm{SL}(4)$, we use $\mathrm{SL}(4)$ nodes for factor graph optimization. A visualization of the factor graph structure showing different edge types can be seen in Fig. 3. Importantly, given that there is error in VGGT’s estimate of calibration, we do not know what the true camera calibration is for the overlapping frames — only that they must have the same calibration.
Intra submap frame alignment. For two keyframes $i$ and $j$ inside a submap, the homography $\mathbf{H}^{i}_{j}$ from Eq. 2 that transfers correspondences between point clouds $\mathbf{X}_{i}$ and $\mathbf{X}_{j}$ is simply a rotation and translation which we can get directly from the VGGT outputted poses as:
@@
Inter submap frame alignment. Now, we have to find the alignment between the overlapping frames $i$ and $j$ in overlapping submaps. For example, nodes $H_{3}$ and $H_{4}$ in Fig. 3 correspond to the same keyframe processed in the first and second submap. Importantly, instead of solving for a 15-DoF transformation by optimizing an alignment over the 3D points corresponding to the two overlapping frames as was done in VGGT-SLAM, here we recognize that the overlapping frames by construction must have the same translation, rotation, and camera calibration. Thus, we enforce that the relative translation and rotation between the overlapping frames is zero and determine the transformation needed to align their calibration. We then solve for the scale factor, $s$.
The homography in Eq. 2 then simplifies to:
@@
where the calibration components, $\textbf{K}_{i}$ and $\textbf{K}_{j}$, of the homography come from the VGGT estimated calibrations. Importantly, even if the calibrations $\textbf{K}_{i}$ and $\textbf{K}_{j}$ do not represent the true camera calibration (which can occur if VGGT incorrectly estimates camera calibration for the current submap), this enforces their calibrations to be identical. We know they must be identical because the overlapping frames represent the same camera. Solving for the scale factor is the only time we directly use the VGGT 3D points. To find $s$, we leverage two ideas. The first is that since the overlapping frames are identical, corresponding 3D points between them are trivially known. The second is that unlike [5], with our new problem formulation, we can easily estimate relative scale separately from other parameters of the homography. The 3D points of each image are defined not w.r.t. the first frame of a submap but rather to their own camera frame. Therefore, corresponding points seen in overlapping keyframes (e.g., $H_{3}$ and $H_{4}$ from Fig. 3) must have the same scale. To estimate the scale, we warp the point clouds $\mathbf{X}_{i}$ and $\mathbf{X}_{j}$ to have the same calibration using $\textbf{K}_{i}^{-1}$ and $\textbf{K}_{j}$ and then estimate the scale factor as the median ratio between corresponding 3D point distances for all points with sufficient depth confidence $\mathcal{C}$.
@@ IV-B Image retrieval verification with VGGT
In this section, we demonstrate that by studying the attention layers of VGGT, we can easily extract information to determine if VGGT is able to find overlap between input images. While we find our analysis holds for an arbitrary number of frames passed to VGGT, since our objective in this section is image retrieval verification, we focus on the scenario where VGGT is given two images (a query and a retrieved frame). In particular, we reveal that layer 22 in VGGT consistently pinpoints whether VGGT predicts correspondences between the two images. In Fig. 4 we visualize the attention matrix of a selected key token (shown with a black star), with respect to all other query tokens. Layer 22 consistently shows a spotlight-like attention map which shows larger attention values at the correspond location of the image. This is not clearly present in the other layers, including the nearby layers (21 and 23) which are shown in Fig. 4 for comparison. This spotlight affects also works on low texture regions of a scene as demonstrated by the first example of Fig. 4 which places the query token on a plain white wall.
@@
(a) Queried and retrieved frames that have overlap. Our estimated match score: 1.026.
(b) Queried and retrieved frames that do not have overlap. Our estimated match score: 0.491.
(c) Queried and retrieved frames that do not have overlap, but SALAD incorrectly estimated a match between them. Our estimated match score: 0.550.